Is y+P(x)y=Q(x)yn a linear equation for integral value of n. Justify your answer.


1
Expert's answer
2021-06-08T19:02:54-0400

SOLUTION:-

"y' + p\left( x \right)y = q\left( x \right){y^n}\quad (*)"

where p(x) and q(x) are continuous functions on the interval we’re working on and n is a real number.

divide the differential equation by "y^n" to get

"y ^{\u2212n}\n y \n\u2032\n +p(x)y^{1\u2212n}\n =q(x)"

We are now going to use the substitution

"v = =y ^{1-n}"

to convert this into a differential equation in terms of v.

As we’ll see this will lead to a differential equation that we can solve. So

"v^\n\u2032\n =(1\u2212n)y ^\n{\u2212n}"

Now, plugging this as well as our substitution into the differential equation gives,

"\\frac{1}{1\u2212n}\n\n\u200b\n v ^\n\u2032\n +p(x)v=q(x)"

This is a linear differential equation that we can solve for v

and once we have this in hand we can also get the solution to the

original differential equation by plugging v.

back into our substitution and solving for y

So the equation is reduced to a linear equation.


If n=0 equation (*) is a linear equation, in our case


n=2 equation is not linear equation, but reduced to a linear equation.